7. A 68.5-kg astronaut is doing a repair in space on the orbiting space

station. She throws a 2.25-kg tool away from her at 3.20 m/s relative to

the space station. With what speed and in what direction will she begin

to move?


1
Expert's answer
2021-10-10T16:00:38-0400

Apply momentum conservation to get the answer:


"0=m_Av_a-m_tv_t,\\\space\\
v_a=\frac{m_tv_t}{m_a}=\frac{2.25·3.2}{68.5}=0.105\text{ m/s}."